OpenStudy (anonymous):

Which of the following is a polynomial with roots 4, −5, and 7?

OpenStudy (anonymous):

A) f(x)=x^3-6x^2-27x+140 B) f(x)=x^3-6x^2-20x+27 C) f(x)=x^3-20x^2-27x+35 D) f(x)=x^3-20x^2-35x+140

OpenStudy (anonymous):

There's a simple way to do this, subsitute the roots for x If it works out to be zero then it is a root If it isn't zero it isn't a root

OpenStudy (anonymous):

uhh yeah i dont get it

OpenStudy (anonymous):

So I'll work out an example of what I'm doing for you Let's try option A first Substituting 4 for A gets you \[4^3-6(4^2)-27(4)+140\] Solving that gets you 0 So 4 is a root of this function, but that's not it you have to test -5 and 7 too And IF it doesn't work for either one of them then A isn't the right answer